a) A total charge of 72 Coulombs is transmitted by a silver wire 1.5 mm in diameter over a time of 1 hour and 30 minutes. What is the current and its density in the wire? 

 



1
Expert's answer
2021-11-04T10:18:14-0400

The current:


"I=Q/t=72/(1.5·3600)=0.013\text{ A}."


Current density:


"J=I/A=4I/(\pi d^2)=4·0.013/(\pi·1.5·10^{-3})=\\=7356\text{ A/m}^2."